# SPARSE: Scattering Poles and Amplitudes from Radial Schrödinger Equations

### v1

@@ 5 Scattering amplitudes and resonances
With this definition, a $T$-matrix element $T_{ij}$ corresponds to the partial-wave scattering amplitude $f_{i\leftarrow j}$ from an initial channel $j$ to a final channel $i$ times the relative momentum in the final channel,
@@
The reactance matrix $K$ is defined as the Caley transform of the $S$-matrix,
@@
(see Section 14-e of Reference [11]). We see from Equations (53) and (54) that unitarity ($S^{\dagger}=S^{-1}$) and time-reversal symmetry ($S^{T}=S$) require that the $K$-matrix is real and symmetric. The $K$-matrix calculated using Equations (44), (48), and (49) is real by definition. It is only approximately symmetric within numerical errors. Let us call this approximately symmetric matrix $\tilde{K}$. SPARSE calculates the degree of asymmetry as
@@
and warns the user if it is larger than a tolerance (1% by default). Then, SPARSE calculates the exactly symmetric $K$-matrix as
$K=\frac{\tilde{K}+\tilde{K}^{t}}{2}.$ (56)
@@
an using (57). The corresponding partial-wave cross-section, $\sigma_{i\leftarrow j}$, is proportional to the square modulus of the scattering amplitudes,
$\sigma_{i\leftarrow j}=\frac{4\pi(2J+1)}{(2J_{\alpha_{j}}+1)(2J_{\beta_{j}}+1)}\lvert f_{i\leftarrow j}\rvert^{2},$ (58)
where $J$ is the total angular momentum and $J_{\alpha_{j}}$ and $J_{\beta_{j}}$ are the spins of the two scattering particles $\alpha_{j}$ and $\beta_{j}$ in the initial channel $j$. The total unpolarized cross section can be obtained by summing over the spins $s_{i}$, $s_{j}$, the orbital angular momenta $l_{j}$, $l_{i}$, and the total angular momentum $J$.
From Equation (57), we see that the scattering amplitude has a pole whenever the denominator on the right side is a singular (i.e., noninvertible) matrix. We can therefore find scattering resonances from the zeros of the function
$F(E)=\det(I-iK(E))$ (59)
for values of $E$ in the complex plane, where we have marked the energy dependence explicitly.
SPARSE calculates the $K$-matrix only for real energies, therefore the zeros of $F(E)$ must be found by extrapolating from the real axis to the complex plane. One way to do this is through the use of a rational approximation, where the function $F(E)$ is approximated by the ratio of two polynomials $N(E)$ and $D(E)$,
$F(E)\approx\frac{N(E)}{D(E)},$ (60)
which allows for the extrapolation to complex values of $E$. SPARSE uses the AAA algorithm for rational polynomial approximation [12] or, more precisely, its implementation as the function AAA in the interpolation module of SciPy [8]. The zeros and poles of $F(E)$ are then found as zeros of $N(E)$ and $D(E)$, respectively.
One known complication when dealing with rational approximations is the occurrence of near-zero residue poles or pole pairs with opposite residues that are so close as to nearly cancel each other. Since the AAA rational approximation typically has as many zeros as poles, the appearance of spurious poles of $F(E)$ would result in spurious zeros, which could be mistakenly identified as scattering resonances. To avoid the occurrence of such spurious poles, the user must first identify them and then adjust the parameters of the AAA algorithm to prevent their appearance. Luckily, it is very easy to identify spurious poles in the rational approximation of $F(E)$. In fact, the poles of $F(E)$ from Equation (59) must coincide with the poles of the $K$-matrix, which are located exclusively on the real energy axis. Since SPARSE can calculate the $K$-matrix for real energies, it is usually possible to roughly establish the number and the positions of the $K$-matrix poles quite reliably. Therefore, the user has only to adjust the parameters of the AAA algorithm until the poles of the rational approximation of $F(E)$ are compatible with the poles of the $K$-matrix. Once the poles have been identified as genuine, the zeros can be reliably identified with scattering resonances.
@@ 6 Showcase
Table 1: Integer label $i$, orbital angular momentum $l$, threshold $T$, and reduced mass $\mu$ for the channels in the SPARSE showcase example. Units are arbitrary.
$i$ $l$ $T$ $\mu$
1 1 0 1000
2 0 100 1000
[…]
@@
$V_{11}(r)$ is a potential well with a short-distance repulsion,
$V_{11}(r)=\biggl{(}\frac{0.05}{r}-100\biggr{)}\exp\bigl{(}-(100r)^{2}\bigr{)}$ (62)
$V_{22}(r)$ is the same potential as $V_{11}(r)$ but shifted by $+100$,
$V_{22}(r)=V_{11}(r)+100$ (63)
and $V_{12}(r)=V_{21}(r)$ is Gaussian coupling potential,
@@
The coordinate space of $r$ is a grid of points in the interval $[0,1]$ with a discretization step $d=10^{-6}$. The potential $V_{11}(r)$ has been engineered so that it has a $P$-wave bound state ($l_{1}=1$) with energy somewhat below $T_{1}=0$. The potential $V_{22}(r)$ has been engineered so that in isolation it would have an $S$-wave bound state ($l_{2}=0$) with energy somewhat below $T_{2}=100$. The coupling potential $V_{12}(r)$ has been engineered so that the would-be bound state in the second channel appears instead as a narrow scattering resonance in the first channel. These potential matrix elements are plotted in Figure 1.
@@
We use SPARSE to calculate the $K$-matrix for a grid of energies in the interval $[0,80]$ with a constant spacing of 0.1. The $K$-matrix has only one element, since there is only one open channel in the given energy region. We plot the $K$-matrix as a function of the energy in Figure 2. We also calculate the square modulus of the $T$-matrix element and plot it in Figure 3.
Figure 2: $K$-matrix for the SPARSE showcase example. The horizontal axis has arbitrary units.
Figure 3: Square modulus of the $T$-matrix for the SPARSE showcase example. The horizontal axis has arbitrary units.
The scattering amplitude looks complicated because it results from the overlap of two different line shapes. The smooth enhancement is due to the bound state in channel 1, which has an energy of -12.5. The sharp oscillation of the amplitude in between 60 and 80 is due to a resonance. That oscillation in $\lvert T\rvert^{2}$ matches a pole of the $K$-matrix at a similar energy. Using the AAA rational approximation to extrapolate from real energies to the complex plane, we find that there is a pole of the $T$-matrix located at $Z=67.02-0.02i$. It corresponds to a resonance with a mass of $M=\Re(Z)=67.02$ and a width of $\Gamma=-2\Im(Z)=0.04$.
The SPARSE repository contains a Python script that calculates the CSV files channels.csv and potential.csv used for this example. It also contains a handy IPython notebook which reproduces the results of this section using the SPARSE Application Programming Interface.
@@ 7 Summary
SPARSE is a simple Python algorithm for the calculation of scattering amplitudes and resonances using the Schrödinger equation. It solves a system of coupled radial Schrödinger equations that represents a partial wave of a two-body scattering process. It reduces the differential equation to a nonhomogeneous linear system by approximating the second derivative using the finite-difference method and imposing Dirichlet boundary conditions. It calculates the numerical wavefunctions by solving the system, then compares them to the analytical wavefunctions of scattering states to obtain the $K$-matrix elements. It calculates the scattering amplitudes from the $K$-matrix elements for real energies, then calculates resonances by extrapolating to complex energies using a rational approximation.
The only inputs required by SPARSE are the orbital angular momenta, reduced masses, and thresholds for the channels of the Schrödinger equation plus the numerical coordinate space and potential matrix. These numerical inputs are provided in the form of two CSV files named channels.csv and potential.csv. The user can calculate these numerical inputs using any computational software. The Application Programming Interface is designed so that it can be used with minimal Python literacy. The SPARSE repository also contains a Python script that creates working examples for the input CSV files. It also contains an IPython notebook which showcases the program’s functionality and serves as a useful tutorial.
@@ Acknowledgements
Appendix A Energy limits
There are limits to which scattering energies are accessible to SPARSE. Here we explain how these limits are calculated from the numerical inputs.
An overall lower energy limit is given by
$E\geq\min_{i}\{T_{i}\},$ (65)
since there can obviously be no scattering for energies below the lowest threshold.
An upper bound on the energy from channels with a finite threshold $T_{i}$ is given by Shannon’s theorem. A wavefunction is completely determined by its values at a grid of evenly spaced points only if its wavelength is greater than twice the discretization step $d$,
$\lambda>2d.$ (66)
Since the momentum is related to the wavelength by
$p=\frac{2\pi}{\lambda},$ (67)
the upper bound on the relative momentum in SPARSE is
$p<\frac{\pi}{d}.$ (68)
Therefore, the overall limit for the energy is
$E<\frac{\pi^{2}}{2\mu_{i}d^{2}}+T_{i}$ (69)
for all $i$ with a finite threshold $T_{i}$. Another upper bound on the energy from confining channels with $T_{i}=\infty$ is given by the corresponding diagonal elements of the potential matrix. Since SPARSE assumes that the wavefunction components in confining channels is exactly zero at the maximum distance $R$, the energy must be much smaller than the corresponding potential at $r=R$,
$E\ll V_{ii}(R)$ (70)
for all $i$ with $T_{i}=\infty$.
The scattering energies cannot be too close to a threshold $T_{i}$, either. There is an upper bound for energies below the threshold, $E<T_{i}$. For a physical solution, the asymptotic wavefunction component in channel $i$ at large $r$ is proportional to
$\exp\bigl{(}-\sqrt{2\mu_{i}(T_{i}-E)}r\bigr{)}.$ (71)
SPARSE assumes that the wavefunction component for channels with $T_{i}>E$ is exactly zero at the maximum distance $r=R$, which is only a good approximation as long as the argument of the exponential satisfies
$\sqrt{2\mu_{i}(T_{i}-E)}R\gg 1.$ (72)
Therefore we obtain an upper bound for energies below the threshold $T_{i}$,
$E-T_{i}\ll-\frac{1}{2\mu_{i}R^{2}}.$ (73)
There is also a lower bound for energies above the threshold, $E>T_{i}$. SPARSE uses Equations (44), (48), and (49) to calculate the $K$-matrix from the numerical wavefunctions of scattering states. For Equations (48) and (49) to be applicable, however, the lower limit of the integrals, $a=R-\pi/p_{i}$, must exceed the radius of the potential, $r_{\Lambda}$, which requires
$E-T_{i}>\frac{\pi^{2}}{2\mu_{i}(R-r_{\Lambda})^{2}}$ (74)
For channels with $l_{i}>0$, a more stringent limit may be obtained from the condition that the argument $p_{i}r$ of the Riccati-Bessel functions is much greater than $l_{i}$ for values in the integration domain of Equations (48) and (49). This requires
$E-T_{i}\gg\frac{(l+\pi)^{2}}{2\mu_{i}R^{2}}.$ (75)

### v2

@@ 5 Scattering poles and amplitudes
Within this convention, a $T$-matrix element $T_{ij}$ corresponds to the partial-wave scattering amplitude $f_{i\leftarrow j}$ from an initial channel $j$ to a final channel $i$ times the relative momentum in the final channel,
@@
The reactance matrix $K$ is defined as the Cayley transform of the $S$-matrix,
@@
(see Section 14-e of Reference [12]). We see from Equations (54) and (55) that unitarity ($S^{\dagger}=S^{-1}$) and time-reversal symmetry ($S^{\mathsf{T}}=S$) require that the $K$-matrix is real and symmetric. The $K$-matrix calculated using Equations (44), (46), and (51) is real by definition. It is only approximately symmetric within numerical errors. Let us call this approximately symmetric matrix $\tilde{K}$. SPARSE calculates the degree of asymmetry as
@@
and warns the user if it is larger than a tolerance (1% by default). Then SPARSE calculates the exactly symmetric $K$-matrix as
$K=\frac{\tilde{K}+\tilde{K}^{\mathsf{T}}}{2}.$ (57)
Resonances may appear as poles of the $K$-matrix on the real axis of the physical Riemann sheet. For energies near the pole, the $K$-matrix elements can be expressed as
$K_{ij}=g_{i}g_{j}\frac{\Gamma/2}{M-E}+K^{\text{nonres}}_{ij}.$ (58)
where $K^{\text{nonres}}_{ij}$ is a “nonresonant” contribution that is regular at the pole energy [^5] Contributions from other poles at different energies may be contained in this term. and $M$, $\Gamma$, and $g_{i}$ are respectively the mass of the resonance, its width, and its coupling to the channel $i$. The couplings $g_{i}$ are normalized, $\sum_{i}g_{i}^{2}=1$, so that the couplings squared give the branching ratios of the resonance,
$B_{i}=g_{i}^{2}.$ (59)
SPARSE calculates scattering poles by interpolating between numerical $K$-matrix elements obtained at a set of discrete energies. The interpolation is performed using the AAA algorithm for rational polynomial approximation [13], which yields the pole energies $M$ and the residues of the $K$-matrix elements at the poles,
$R_{ij}=\operatorname{Res}_{M}(K_{ij}).$ (60)
Not all poles of the $K$-matrix correspond to resonances. SPARSE identifies $K$-matrix poles associated to resonances by comparing the residues calculated from the AAA algorithm with the analytic pole term in Equation (58). If the $K$-matrix pole indeed corresponds to a resonance, then Equations (58) and (60) require
$R_{ij}=-g_{i}g_{j}\Gamma/2,$ (61)
which implies two constraints on the residues:
$R_{ii}<0\qquad\text{and}\qquad R_{ij}^{2}=R_{ii}R_{jj}$ (62)
for every $i$ and $j$. SPARSE calculates the residues $R_{ij}$ and checks that they satisfy the constraints in Equation (62). If the constraints are satisfied, SPARSE calculates the width of the resonance as
$\Gamma=-2\sum_{i}R_{ii},$ (63)
then it calculates the couplings as
$g_{i}=\operatorname{sgn}(-R_{1i})\sqrt{-2R_{ii}/\Gamma},$ (64)
where the sign function on the right side is necessary to fix the relative signs between the couplings so that they agree with Equation (61). Note that the couplings are defined up to an overall sign, that is, the $K$-matrix in Equation (58) is invariant if we substitute $g_{i}\to-g_{i}$ for every $i$. In writing Equation (64), we have chosen the convention $g_{i}\geq 0$ for $i=1$.
@@
and using Equation (53). The corresponding partial-wave cross-section, $\sigma_{i\leftarrow j}$, is proportional to the square modulus of the scattering amplitudes,
$\sigma_{i\leftarrow j}=\frac{4\pi(2J+1)}{(2J_{1,j}+1)(2J_{2,j}+1)}\lvert f_{i\leftarrow j}\rvert^{2},$ (66)
where $J$ is the total angular momentum and $J_{1,j}$ and $J_{2,j}$ are the spins of the particles 1 and 2 in the initial channel $j$. The total unpolarized cross section can be obtained by summing over the spins $s_{i}$, $s_{j}$, the orbital angular momenta $l_{i}$, $l_{j}$, and the total angular momentum $J$.
It is worth noting that the mass $M$, width $\Gamma$, and couplings $g_{i}$ obtained from the $K$-matrix pole are generally not the same as the parameters of the bona fide resonance, that is, the $T$-matrix pole. However, the parameters of the $K$-matrix and $T$-matrix poles do coincide in the case of an isolated resonance. To see this, insert Equation (58) into Equation (65) keeping only the pole term and observe that the result is the resonant $T$-matrix element:
$T_{ij}\approx ig_{i}g_{j}\frac{i\Gamma/2}{E-M+i\Gamma/2}.$ (67)
If the resonance is not isolated, then the parameters $M$, $\Gamma$, and $g_{i}$ obtained from the $K$-matrix pole can differ significantly from those of the $T$-matrix pole. Nonetheless, the $K$-matrix pole parameters are physically relevant as they can be related to $K$-matrix fits to experimental data. In those cases where a distinction is necessary, $M$, $\Gamma$, and $g_{i}$ may be referred to as the nominal mass, width, and couplings of the resonance.
@@ 6 Showcase
Table 1: Integer label $i$, orbital angular momentum $l$, reduced mass $\mu$, and threshold $V^{\infty}$ for the channels in the SPARSE showcase example. Units are arbitrary.
$i$ $l$ $\mu$ $V^{\infty}$
1 1 1000 0
2 0 1000 100
Next, we examine an application of the SPARSE algorithm to a relatively simple system with two coupled channels. Units are arbitrary. The two channels are described in Table 1. The potential matrix is
@@
The first diagonal element $V_{11}(r)$ is a potential well with a short-distance repulsion,
$V_{11}(r)=\biggl(\frac{0.05}{r}-100\biggr)\exp\bigl(-(100r)^{2}\bigr).$ (69)
The second diagonal element $V_{22}(r)$ is the same potential well but shifted by +100+100,
$V_{22}(r)=V_{11}(r)+100.$ (70)
The offdiagonal element $V_{12}(r)=V_{21}(r)$ is Gaussian coupling potential,
@@
The coordinate space of $r$ is a grid of points in the interval $[0,1]$ with a discretization step $d=10^{-6}$. The potential $V_{11}(r)$ has been engineered so that it has a $P$-wave (that is, $l=1$) bound state with energy somewhat below the first threshold, $V^{\infty}_{1}=0$. The potential $V_{22}(r)$ has been engineered so that in isolation it would have an $S$-wave (that is, $l=0$) bound state with energy somewhat below the second threshold, $V^{\infty}_{2}=100$. The coupling potential $V_{12}(r)$ has been engineered so that the would-be bound state in the second channel appears instead as a narrow scattering resonance in the first channel. These potential matrix elements are plotted in Figure 1.
@@
We use SPARSE to calculate the $K$-matrix for a grid of energies in the interval $[0,80]$ with a constant spacing of 0.05. The $K$-matrix has only one element, since there is only one open channel in this energy region. We plot the $K$-matrix as a function of the energy in the left panel of Figure 2. We also calculate the square modulus of the $T$-matrix element and plot it in the right panel of Figure 2.
Figure 2: $K$-matrix (left) and square modulus of the $T$-matrix (right) in the SPARSE showcase example. The horizontal axes have arbitrary units.
The scattering amplitude looks complicated because it results from the overlap of two different line shapes. The smooth enhancement is due to the bound state in channel 1, which has an energy of -12.5. The sharp oscillation of the amplitude in between 60 and 80 is due to a resonance. That oscillation in $\lvert T\rvert^{2}$ matches a pole of the $K$-matrix with mass $M=67.09$ and width $\Gamma=0.66$.
The SPARSE repository contains a Python script that prepares the CSV files channels.csv and potential.csv used for this example. It also contains a handy IPython notebook which reproduces the results of this section using the SPARSE Application Programming Interface. The reader is invited to go through the IPython notebook as a tutorial for using SPARSE.
@@ 7 Summary
SPARSE is a simple and efficient Python algorithm for the calculation of amplitudes and resonances for inelastic scattering in nonrelativistic quantum mechanics. SPARSE solves a system of coupled radial Schrödinger equations that represents a partial wave of a two-body scattering process in systems where the particles can change their state as a result of the collision. The SPARSE algorithm is optimized for the application to systems with tens of coupled Schrödinger equations. It reduces the system of differential equations to a nonhomogeneous linear system by approximating the second derivative using the finite-difference method and imposing Dirichlet boundary conditions. It calculates the numerical wavefunctions by solving the system, then compares them to the analytical wavefunctions of scattering states to obtain the $K$-matrix elements. Finally, SPARSE calculates scattering amplitudes and poles from the $K$-matrix elements for real energies.
The only inputs required by SPARSE are the orbital angular momenta, reduced masses, and thresholds for the each of the Schrödinger equations in the system plus the numerical coordinate space and potential matrix. These numerical inputs are provided by the user in the form of two CSV files named channels.csv and potential.csv. The user can calculate these numerical inputs using any computational software. The SPARSE Application Programming Interface is designed so that it can be used with minimal Python literacy. The SPARSE repository also contains a Python script that creates working examples for the input CSV files. It also contains an IPython notebook which showcases the program’s functionality and serves as a useful tutorial.